# A radial scalar product for Kerr quasinormal modes

### v3

@@ III.4 QNM solutions of Teukolsky’s radial equation
×$\quad\quad\times\xi^{-(iM\tilde{\omega}+s)+\frac{i\left(am-2M^{2}\tilde{\omega}\right)}{2\delta}}\;.$
@@
While we refer to the effect of Eqs. (III.4) and (14) as a similarity transformation, it is sometimes referred to as an elementary power transformation, or a homotopic transformation Ronveaux and Arscott (1995); Chen and Jing (2023); Cho (2020). The reader should note that Eqs. (13a) and (14) contain a coordinate transformation ($r$ to $\xi$), as well as a similarity transformation.
@@
There, the constants, $\text{C}_{0}$ to $\text{C}_{4}$, are defined in Eq. (III.4) using the physical parameters discussed previously.
@@
Equation (15) has a few special properties. We have started with Boyer-Lindquist coordinates, but one also arrives at Eq. (15) if one instead begins with Kerr Ingoing or Kerr Outgoing coordinates Teukolsky and Press (1974). One also arrives at Eq. (15) (albeit with $s\rightarrow-s$ and $A\rightarrow A-2s$) if one uses the alternative tetrad choice discussed in Ref. Teukolsky and Press (1974), which amounts to $t\rightarrow-t$ and $\phi\rightarrow-\phi$. As noted in e.g. Refs. Teukolsky and Press (1974); Whiting (1989), the use of either $s=2$ or $s=-2$ results in mathematically different descriptions of the same gravitational radiation.
@@
Equation (15) has regular singular points at $\xi\in\{-\infty,0\}$ and one irregular singular point at $\xi=1$. Thus Eq. (15) is an instance of the confluent Heun differential operator Ronveaux and Arscott (1995). This is also true in the Schwarzschild limit. While it is sometimes said that Ref. Leaver (1986) treats Eq. (15) as what is called a generalized spheroidal operator, we note here that this is a slight misnomer Ronveaux and Arscott (1995); Hortacsu (2012). Equation (15) may only be transformed into a generalized spheroidal operator if one applies an asymptotic boundary condition at $r=r_{-}$ that is inconsistent with QNMs Leaver (1985); Cook and Zalutskiy (2014). Appendix (C) provides an expanded disambiguation between different forms of the confluent Heun equation and the radial equation.
@@
Equations (13b-III.4) are commonly used to define series solutions for $f(\xi)$. The requirement that these series solutions converge on $\xi\in[0,1)$ happens to be equivalent to finding eigenvalues, $A$, such that certain continued fraction are zero Leaver (1985, 1986); Mano et al. (1996). In particular, see Ref.Leaver (1986) for a proof of convergence for series solutions, example applications of continued fractions to the radial and angular eigen-problems, and how the eigenvalues may be determined by searching for roots of continued fractions.
@@
From the use of continued fractions, values of $A$ are well known to be discrete, and by convention are labeled by the size of their imaginary part: $n=0$ corresponds to the eigenvalue with the smallest value of $|1/\tau|$, and $n=1$ has the next-smallest value Berti et al. (2016); Nollert (1999); Kokkotas and Schmidt (1999). However, the use of continued fractions has arguably obscured deeper understanding of these eigenvalues, and their related solutions. In search of that understanding, it appears that there has been relatively little application of Sturm-Liouville theory to Eq. (15) in the BH context Leaver (1985); Nollert (1999); Berti and Cardoso (2006); Ripley (2022); Lin and Qian (2023); Press (1971); Motl and Neitzke (2003); Ronveaux and Arscott (1995). Nevertheless, since the QNM eigenvalues are well known to be unique for most physical scenarios  [^7] There is some ambiguity in how one defines the eigenvalues. In particular, the fact that $\text{C}_{0}$ is non-zero means that the eigenvalue that one starts with, $A$, need not strictly define the eigenvalue that is used in Eq. (9a). Thus in e.g. the limit that $a=0$, where $A=\ell(\ell+1)-s(s+1)$ for every radial eigenfunction, the effective eigenvalue, $A-\text{C}_{0}$, may be a unique function of $\tilde{\omega}$.  (e.g.  London and Fauchon-Jones (2019); Cook (2020); Seidel (1989); Yang et al. (2012)), the QNM’s themselves are necessarily linearly independent [^8] More precisely, finite sets of QNMs will be linearly independent when their eigenvalues are unique. Infinite sets of QNMs will be minimal if their eigenvalues remain distinct as some order label on the set goes to infinity Christensen (2003); London (2023)., and thus well suited to Sturm-Liouville theory Courant and Hilbert. (1954).
@@ IV Definition of a radial scalar product
The aim of Sturm-Liouville theory is to understand the existence and asymptotic behavior of eigenvalues and eigenfunctions Pinchover and Rubinstein (2005); Arfken et al. (2013); Courant and Hilbert. (1954). Here, we will focus on a facet of the theory that pertains to whether there exists a scalar product and an associated weight function, for which $\mathcal{L}_{\xi}$ is self-adjoint.
@@ IV.1 The scalar product and the weight function
[…]
@@
Since there is no conjugation involved in Eq. (17), the scalar-product shown is not conjugate-symmetric. This feature distinguishes it from an inner-product, and allows analysis of $\mathcal{L}_{\xi}$, irrespective of its complex valued coefficients. The analysis of immediate interest pertains whether $\text{W}(\xi)$ may be defined such that $\mathcal{L}_{\xi}$ is formally self-adjoint Pinchover and Rubinstein (2005); Courant and Hilbert. (1954).
@@
Chapter $5$ Section 3 of Ref. Courant and Hilbert. (1954) provides a detailed overview of how $\mathcal{L}_{\xi}$ may be explicitly defined using integration by parts. In that derivation, the fact that the coefficient of the second derivative term in Eq. (15) vanishes at $\xi=0$ and $\xi=1$ is key to Sturm-Liouville boundary conditions being satisfied Pinchover and Rubinstein (2005). There, it is also shown that $\mathcal{L}_{\xi}$ is formally self-adjoint, i.e.
$\mathcal{L}{{}_{\xi}}\;$ =,$=\;{\mathcal{L}{{}_{\xi}}}^{\dagger}\;,$ (19)
@@
In Eq. (IV.1), $p_{0}(\xi)$ and $p_{2}(\xi)$ are the coefficient functions seen in zero’th and second derivative terms of Eq. (15), respectively. If we define $p_{1}$ to be the coefficient function of $\partial_{\xi}$ in Eq. (15), then
@@
$^{-1}\,\partial_{\xi}\text{W}\;$ W $=\;p_{2}^{-1}(p_{1}-\partial_{\xi}p_{2})\;.$ (22b)
@@
Equation (22b) can be solved for W using the coefficient functions defined in Eq. (IV.1),
[…]
@@
[…]
@@ IV.2 Properties and non-uniqueness of the weight function
With Eq. (26), we arrive at one of this article’s key results. Equation (26) along with Eqs. (III.4) and (IV.1) define a scalar product for which Teukolsky’s radial equation, Eq. (15), is self-adjoint. In turn, self-adjointness, in principle, implies that as long as the eigenvalues may be defined, and that they are unique, then the eigenfunctions are orthogonal, and very likely complete Ronveaux and Arscott (1995). However, there are good reasons why the explicit demonstration of orthogonality and completeness are beyond the scope of this article. The most pressing reason has to do with the fact that $\text{W}(\xi)$, naively, appears to diverge on the event horizon, and at spatial infinity. Understanding that apparent divergence and its consequences is preliminary to further investigation.
Visual inspection of Eq. (IV.1) points to the origin of these divergences. In Eq. (23a), the factor
@@
The first perspective is that the form of the weight function is closely linked to the scaling function, $\mu(\xi)$, defined in Eq. (12). In particular, it is easy show (see Appx. B) that
@@
This weight function is independent of $\tilde{\omega}$, $m$ and $a$, and it is directly relevant for the original radial functions, $R(\xi)$ (see Eq. 9b). In essence, Eq. (30) defines how the weight function changes under similarity transformations on the main problem. One way to arrive at Eq. (30) is to note that such transformations preserve the value of scalar products. While $\text{W}_{0}$ may be of interest for future work, it is not directly considered here because the corresponding differential operator does not support polynomial solutions.
The second perspective is that $\text{W}(\xi)$ is not the only weight function for which $\mathcal{L}_{\xi}$ is self-adjoint. It is left as an exercise for the reader to show that one can actually enforce a strictly positive and real valued weight function,
@@
This exercise follows exactly the steps discussed in Eqs. (IV.1-IV.1), but with the requirement that the weight function is wrapped within an absolute value. While $\mathcal{W}(\xi)$ is amenable to the direct integration approach that we will discuss shortly, it is not at all clear to the author how it may be used with the analytic continuation approach that will be discussed in Sec. V.2. Further, there do not appear to be significant gains in using $\mathcal{W}(\xi)$ rather than $\text{W}(\xi)$. Since complexification of the radial coordinate is also needed for $\mathcal{W}(\xi)$, scalar products will in general still be complex valued. $\mathcal{W}(\xi)$ is noted for potential future interest. Henceforth, this article will focus on the evaluation of scalar products with the comparatively simpler weight function, $\text{W}(\xi)$.
@@
Figure 1: Examples integration paths (white curves) against weight function contours (colored contours) for the $(s,\ell,m,n)=(-2,2,2,3)$ QNM at three BH spins: (left) $a/M=0.7$, $M\tilde{\omega}=0.4713-0.5843i$, (center) $a/M=0.99$, $M\tilde{\omega}=0.8695-0.2058i$, and (right) $a/M=0.999$, $M\tilde{\omega}=0.9558-0.07299i$. Contours for the real and imaginary parts of the weight function, $(\xi)=\mathrm{Re}(\text{W})+i\,\mathrm{Im}(\text{W})$ W ), are shown. For $\mathrm{Re}(\text{W})$, colors between cyan and magenta represent dimensionless values between $-50$ and $50$ respectively. For $\mathrm{Im}(\text{W})$, colors between magenta and yellow represent values between $-50$ and $50$ respectively. Convergence of contours at spatial infinity ($\mathrm{Re}(\xi)=1$) corresponds to the divergence of the weight function. Horizontal contours along $\mathrm{Re}(\xi)<0$ and $\mathrm{Re}(\xi)>0$ are branch cuts. For left, center and right panels, coordinate parameters, $(K_{0},K_{2})$, are $(0.8530,23.4621)$, $(0.7105,1.4010\times 10^{-4})$ and $(0.1585,2.5088\times 10^{-4})$, respectively. Dotted horizontal and vertical lines mark the respective locations of the real and imaginary axes.
@@ V Evaluation of scalar products
[…]
@@ V.1 Direct integration
[…]
@@
$\xi(z)\;=\;\frac{{x_{2}}z^{{x_{0}}}}{({x_{2}}+1)z^{{x_{0}}}-1}\;.$ (34)
In Eq. (34), $x_{0}$ and $x_{2}$ are coordinate parameters that modify $\text{B}_{0}$ and $\text{B}_{2}$, respectively. We will consider $\text{Re}(x0)>0$, such that $\xi(0)=0$. The dependence of Eq. (34) on $x_{2}$ is motivated by the complex radial coordinate used in Ref. Mano et al. (1996). It may also be developed by placing the Laguerre polynomial equation into a $\xi$-like compactified coordinate, and then studying that system’s weight function. The dependence of Eq. (34) on $x_{0}$ is motivated by the observation that $z^{x_{0}}$ is the simplest apparent coordinate modification that impacts singular behavior arising from $\text{B}_{0}$.
@@
[…]
In Eq. (35), we have used the fact that $d\xi=\partial_{z}\xi(z)\,dz$ to write the integral explicitly in $z$. According to Eq. (34), $\xi(z=0)=0$, and $\xi(z=1)=1$; thus, in Eq. (35), the limits of integration do not change.
@@
×$\quad\quad\times\;\left(1-(\text{x}_{2}+1)z^{\text{x}_{0}}\right)^{-\text{B}_{0}-\text{B}_{1}-2}\;.$
While Eq. (34) provides a coordinate that modifies the boundary behavior of $\text{W}(\xi)$, it is Eq. (V.1) that must be analyzed to determine the special values of $x_{0}$ and $x_{2}$ that prevent the effective weight function from diverging at the domain boundaries. In essence, we have two independent conditions: (i) $\text{W}_{\text{eff.}}(0)=0$, and (ii) $\text{W}_{\text{eff.}}(1)=0$. We also have two unknowns: $x_{0}$ and $x_{1}$. Thus, $x_{0}$ and $x_{1}$ may be determined (up to positive constants) by careful analysis of $\text{W}_{\text{eff.}}$.
This analysis is somewhat tedious, but tractable. First, one should focus on the factor of $z^{\text{x}_{0}(\text{B}_{0}+1)-1}$ within $\text{W}_{\text{eff.}}$. If $x_{0}$ is chosen appropriately, then the real part of $\text{x}_{0}(\text{B}_{0}+1)-1$ should equal some positive constant, $\text{K}_{0}$, thereby allowing $\text{W}_{\text{eff.}}$ to go to zero as $z$ goes to zero. Since, $\text{B}_{0}$ is complex valued, $x_{0}$ must be also. These ideas result in the following constraints,
$\text{K}_{0}$ $=-1+\left[1+\text{Re}(\text{B}_{0})\right]\text{Re}(x_{0})-\text{Im}(\text{B}_{0})\text{Im}(x_{0}),$ (37a)
@@
If we assume that $\text{Re}(x_{0})$ only depends on $\text{Re}(\text{B}_{0})$, then then the result for $x_{0}$ follows,
@@
Given, $x_{0}$, one may then determine $x_{2}$. Returning to $\text{W}_{\text{eff.}}$, if its exponential factor has a negative argument as $z\rightarrow 1$, then $\text{W}_{\text{eff.}}$ is assured to be regular in that limit. To that end, one should consider $e^{\frac{\text{B}_{2}\text{x}_{2}}{z^{\text{x}_{0}}-1}}$, and seek $x_{2}$ such that
@@
In Eq. (V.1), argmin outputs the values for which $S$ is minimized in $\text{K}_{0}$ and $\text{K}_{2}$, $\text{mean}_{z}$ computes the unweighted mean over values of $z$, $\text{min}_{z}$ returns the smallest value of a quantity in $z$, and $\text{max}_{z}$ returns the largest value of a quantity in $z$.
In practice, one has a number of remaining options for actually evaluating Eq. (35). For the evaluation of Eq. (V.1), it is found to be useful to densely sample, on log-scale, $z$ near 00 and $1$. The coordinate freedom associated with $x_{0}$ introduces oscillations into $\xi(z)$ which must be well resolved for the accurate determination of $\text{K}_{0}$ and $\text{K}_{2}$. Similarly, for the evaluation of Eq. (35), one will want to densely sample near the boundaries; it is found that $z=\frac{1}{2}(1+\sin(t\pi/2))$, with $t\in(-1,1)$ is sufficient for this purpose. For the determination of $\text{K}_{0}$ and $\text{K}_{2}$, and for the evaluation of Eq. (35), it is found that at least $10^{4}$ points in $z$ are needed for robust results. With these details in hand, it is now possible to compute the integration path, $\xi(z)$, and related scalar-products, Eq. (35), using e.g. $5^{\text{th}}$ order spline integration Virtanen et al. (2020).

### v4

@@ III.4 QNM solutions of Teukolsky’s radial equation
$\quad\quad\times\xi^{-(iM\tilde{\omega}+s)+\frac{i\left(am-2M^{2}\tilde{\omega}\right)}{2\delta}}\;.$
@@
While we refer to the effect of Eqs. (13) and (14) as a similarity transformation, it is sometimes referred to as an elementary power transformation, or a homotopic transformation Ronveaux and Arscott (1995); Chen and Jing (2023); Cho (2020). The reader should note that Eqs. (13a) and (14) contain a coordinate transformation ($r$ to $\xi$), as well as a similarity transformation.
@@
There, the constants, $\text{C}_{0}$ to $\text{C}_{4}$, are defined in Eq. (16) using the physical parameters discussed previously.
@@
Equation (15) has a few special properties. We have started with Boyer-Lindquist coordinates, but one also arrives at Eq. (15) if one instead begins with Kerr Ingoing or Kerr Outgoing coordinates Teukolsky and Press (1974). One also arrives at Eq. (15) (albeit with $s\rightarrow-s$ and $A\rightarrow A-2s$) if one uses the alternative tetrad choice discussed in Ref. Teukolsky and Press (1974), which amounts to $t\rightarrow-t$ and $\phi\rightarrow-\phi$. As noted in e.g. Refs. Teukolsky and Press (1974); Whiting (1989), the use of either $s=2$ or s=2s=-2 results in mathematically different descriptions of the same gravitational radiation.
@@
Equation (15) has regular singular points at $\xi\in\{-\infty,0\}$ and one irregular singular point at $\xi=1$. Thus Eq. (15) is an instance of the confluent Heun differential operator Ronveaux and Arscott (1995). This is also true in the Schwarzschild limit. While it is sometimes said that Ref. Leaver (1986) treats Eq. (15) as what is called a generalized spheroidal operator, we note here that this is a slight misnomer Ronveaux and Arscott (1995); Zarrinkamar and Hortaçsu (2018). Equation (15) may only be transformed into a generalized spheroidal operator if one applies an asymptotic boundary condition at $r=r_{-}$ that is inconsistent with QNMs Leaver (1985); Cook and Zalutskiy (2014). Appendix (C) provides an expanded disambiguation between different forms of the confluent Heun equation and the radial equation.
@@
Equations (13b-16) are commonly used to define series solutions for $f(\xi)$. The requirement that these series solutions converge on $\xi\in[0,1)$ happens to be equivalent to finding eigenvalues, $A$, such that certain continued fractions are zero Leaver (1985, 1986); Mano et al. (1996). In particular, see Ref.Leaver (1986) for a proof of convergence for series solutions, example applications of continued fractions to the radial and angular eigen-problems, and how the eigenvalues may be determined by searching for roots of continued fractions.
@@
From the use of continued fractions, values of $A$ are well known to be discrete. However, the use of continued fractions has arguably obscured deeper understanding of the distribution of eigenvalues, $A$, and their related radial (labeled in $n$) and angular functions (labeled in $\ell$ and $m$). In search of that understanding, it appears that there has been relatively little application of Sturm-Liouville theory to Eq. (15) in the BH context Leaver (1985); Nollert (1999); Berti and Cardoso (2006); Ripley (2022); Lin and Qian (2023); Press (1971); Motl and Neitzke (2003); Ronveaux and Arscott (1995); Zimmerman et al. (2015). Nevertheless, since the Kerr QNMs’ eigenvalues are well known to be unique for most physical scenarios  [^7] There is some ambiguity in how one defines the eigenvalues. In particular, the fact that $\text{C}_{0}$ is non-zero means that the eigenvalue that one starts with, $A$, need not strictly define the eigenvalue that is used in Eq. (9a). Thus in e.g. the limit that $a=0$, where $A=\ell(\ell+1)-s(s+1)$ for every radial eigenfunction, the effective eigenvalue, $A-\text{C}_{0}$, may be a unique function of $\tilde{\omega}$.  (e.g.  London and Fauchon-Jones (2019); Cook (2020); Seidel (1989); Yang et al. (2012)), the QNM’s themselves are necessarily linearly independent [^8] More precisely, finite sets of QNMs will be linearly independent when their eigenvalues are unique. Infinite sets of QNMs will be minimal if their eigenvalues remain distinct as some order label on the set goes to infinity Christensen (2003); London (2023)., and thus well suited to Sturm-Liouville theory Courant and Hilbert (1989).
@@ IV Definition of a radial scalar product
The aim of Sturm-Liouville theory is to understand the existence and asymptotic behavior of eigenvalues and eigenfunctions Pinchover and Rubinstein (2005); Arfken et al. (2013); Courant and Hilbert (1989). Here, we will focus on a facet of the theory that pertains to whether there exists a scalar product and an associated weight function, for which $\mathcal{L}_{\xi}$ is self-adjoint. While a scalar product may be developed under which the problem is self-adjoint (symmetric under scalar products), it is still the case the the QNM radial problem is non-Hermitian; e.g. it has non-trivially complex eigenvalues.
The immediate aim of what follows is to derive a scalar product and related weight function such that Eq. (15) is formally self-adjoint. In other words, a function, $\text{W}(\xi)$, is determined such that $\text{W}(\xi)\mathcal{L}_{\xi}$ is in Sturm-Liouville form. As the scalar product will ultimately be an integral over $\xi$ involving $\text{W}(\xi)$, the behavior $\text{W}(\xi)$ at the domain boundaries is addressed in Sec. V.
@@ IV.1 The scalar product and the weight function
${\langle{\mathrm{a}}\,|\,{\mathrm{b}}\rangle}\;=\;\int_{0}^{1}\,\mathrm{a}(\xi)\mathrm{b}(\xi)\;\text{W}(\xi)\;d\xi\;.$ (17)
@@
Since there is no conjugation involved in Eq. (17), the scalar-product shown is not conjugate-symmetric. This feature distinguishes it from an inner-product, and allows analysis of $\mathcal{L}_{\xi}$, irrespective of its complex valued coefficients. The analysis of immediate interest pertains whether $\text{W}(\xi)$ may be defined such that $\mathcal{L}_{\xi}$ is formally self-adjoint Pinchover and Rubinstein (2005); Courant and Hilbert (1989).
@@
Chapter $5$ Section 3 of Ref. Courant and Hilbert (1989) provides a detailed overview of how $\mathcal{L}_{\xi}$$^{\dagger}$ may be explicitly defined using integration by parts. In that derivation, the fact that the coefficient of the second derivative term in Eq. (15) vanishes at $\xi=0$ and $\xi=1$ is key to Sturm-Liouville boundary conditions being satisfied Pinchover and Rubinstein (2005). There, it is also shown that $\mathcal{L}_{\xi}$ is formally self-adjoint, i.e.
$\mathcal{L}{{}_{\xi}}\;$ $=\;{\mathcal{L}{{}_{\xi}}}^{\dagger}\;$, (19)
@@
In Eq. (20), $p_{0}(\xi)$ and $p_{2}(\xi)$ are the coefficient functions seen in zero’th and second derivative terms of Eq. (15), respectively. If we define $p_{1}$ to be the coefficient function of $\partial_{\xi}$ in Eq. (15), then
@@
$\text{W}^{-1}\,\partial_{\xi}\text{W}\;$ $=\;p_{2}^{-1}(p_{1}-\partial_{\xi}p_{2})\;.$ (22b)
@@
Equation (22b) can be solved for W using the coefficient functions defined in Eq. (21),
$\text{W}(\xi)\;$ $=\;\,\frac{1}{p_{2}(\xi)}\,\exp\left(\int^{\xi}\frac{p_{1}(\xi^{\prime})}{p_{2}(\xi^{\prime})}\,d\xi^{\prime}\,\right)$ (23a)
@@
${\langle{\mathrm{a}}\,|\,{\mathrm{b}}\rangle}\;=\;\int_{0}^{1}\,\mathrm{a}(\xi)\,\mathrm{b}(\xi)\;\;\xi^{\text{B}_{0}}(1-\xi)^{\text{B}_{1}}e^{\frac{\text{B}_{2}}{1-\xi}}\;d\xi\;.$ (26)
@@ IV.2 Properties and non-uniqueness of the weight function
With Eq. (26), we arrive at one of this article’s key results. Equation (26) along with Eqs. (16) and (24) define a scalar product for which Teukolsky’s radial equation, Eq. (15), is self-adjoint. In turn, self-adjointness, in principle, implies that as long as the eigenvalues may be defined, and that they are unique, then the eigenfunctions are orthogonal, and very likely complete Ronveaux and Arscott (1995). However, there are good reasons why the explicit demonstration of orthogonality and completeness are beyond the scope of this article [^10] In particular, one route to demonstrating completeness of the single-frequency radial functions would be to construct an isomorphism between them, and a known basis London (2023); Christensen (2003).. The most pressing reason has to do with the fact that $\text{W}(\xi)$, naively, appears to diverge on the event horizon, and at spatial infinity. Understanding that apparent divergence and its consequences is preliminary to further investigation.
Visual inspection of Eq. (23) points to the origin of these divergences. In Eq. (23a), the factor
@@
The first perspective is that the form of the weight function is closely linked to the scaling function, $\mu(\xi)$, defined in Eq. (12). In particular, it is easy show (see e.g. Eq. 145) that
@@
This weight function is independent of $\tilde{\omega}$, $m$ and $a$, and it is directly relevant for the original radial functions, $R(\xi)$ (see Eq. 9b). In essence, Eq. (30) defines how the weight function changes under similarity transformations on the main problem. One way to arrive at Eq. (30) is to note that such transformations preserve the value of scalar products. While $\text{W}_{0}$ has already been used in operator perturbation theory (e.g. Zimmerman et al. (2015)), it is not directly considered here because the corresponding differential operator does not support polynomial solutions that are compatible with the QNM boundary conditions.
The second perspective is that $\text{W}(\xi)$ is not the only weight function for which $\mathcal{L}_{\xi}$ is formally self-adjoint. It is left as an exercise for the reader to show that one can actually enforce a strictly positive and real valued weight function,
@@
This exercise follows exactly the steps discussed in Eqs. (20-23), but with the requirement that the weight function is wrapped within an absolute value. Further, one must use the fact that $\partial_{\xi}\mathcal{W}=\text{sign}(\text{W})\,\partial_{\xi}\text{W}$. While $\mathcal{W}(\xi)$ is amenable to the direct integration approach that we will discuss shortly, it is not at all clear to the author how it may be used with the analytic continuation approach that will be discussed in Sec. V.2. Further, there do not appear to be significant gains in using $\mathcal{W}(\xi)$ rather than $\text{W}(\xi)$. Since complexification of the radial coordinate is also needed for $\mathcal{W}(\xi)$, scalar products will in general still be complex valued. $\mathcal{W}(\xi)$ is noted for potential future interest. Henceforth, this article will focus on the evaluation of scalar products with the comparatively simpler weight function, $\text{W}(\xi)$.
@@
Figure 1: Examples integration paths (white curves) against weight function contours (colored contours) for the $(s,\ell,m,n)=(-2,2,2,3)$ QNM at three BH spins: (left) $a/M=0.7$, $M\tilde{\omega}=0.4713-0.5843i$, (center) $a/M=0.99$, $M\tilde{\omega}=0.8695-0.2058i$, and (right) $a/M=0.999$, $M\tilde{\omega}=0.9558-0.07299i$. Contours for the real and imaginary parts of the weight function, $\text{W}(\xi)=\mathrm{Re}(\text{W})+i\,\mathrm{Im}(\text{W})$, are shown. For $\mathrm{Re}(\text{W})$, colors between cyan and magenta represent dimensionless values between 50-50 and $50$ respectively. For $\mathrm{Im}(\text{W})$, colors between magenta and yellow represent values between 50-50 and $50$ respectively. Convergence of contours at spatial infinity ($\mathrm{Re}(\xi)=1$) corresponds to the divergence of the weight function. Horizontal contours along $\mathrm{Re}(\xi)<0$ and $\mathrm{Re}(\xi)>0$ are branch cuts. For left, center and right panels, coordinate parameters, $(K_{0},K_{2})$, are $(0.8530,23.4621)$, $(0.7105,1.4010\times 10^{-4})$ and $(0.1585,2.5088\times 10^{-4})$, respectively. Dotted horizontal and vertical lines mark the respective locations of the real and imaginary axes.
@@ V Evaluation of scalar products
${\langle{\text{a}}\,|\,{\text{b}}\rangle}\;=\;\int_{0}^{1}\,\text{a}(\xi)\,\text{b}(\xi)\;\;\xi^{\text{B}_{0}}(1-\xi)^{\text{B}_{1}}e^{\frac{\text{B}_{2}}{1-\xi}}\;d\xi\;.$ (33)
@@ V.1 Direct integration
An extremely useful lesson from calculus is that, given an integral, we may change the coordinate representation of its integrand without changing the value of the integral. So while the integrand of Eq. (33) is coordinate dependent, its value, $\langle{\text{a}}\,|\,{\text{b}}\rangle$}, is coordinate independent. This concept leads us to seek a new real valued coordinate, $z$, such that the integrand of Eq. (33) does not diverge. Henceforth, we will use $\xi(z)$ to refer to a complex valued integration path, and $\xi$ to refer to the quantity defined by Eq. (10).
@@
$\xi(z)\;=\;\frac{{\text{x}_{2}}z^{{\text{x}_{0}}}}{({\text{x}_{2}}+1)z^{{\text{x}_{0}}}-1}\;.$ (34)
In Eq. (34), $\text{x}_{0}$ and $\text{x}_{2}$ are coordinate parameters that modify $\text{B}_{0}$ and $\text{B}_{2}$, respectively. We will consider $\text{Re}(x0)>0$, such that $\xi(0)=0$. The dependence of Eq. (34) on $\text{x}_{2}$ is motivated by the complex radial coordinate used in Ref. Mano et al. (1996). It may also be developed by placing the Laguerre polynomial equation into a $\xi$-like compactified coordinate, and then studying that system’s weight function. The dependence of Eq. (34) on $\text{x}_{0}$ is motivated by the observation that $z^{\text{x}_{0}}$ is the simplest apparent coordinate modification that impacts singular behavior arising from $\text{B}_{0}$.
@@
${\langle{\text{a}}\,|\,{\text{b}}\rangle}=\int_{0}^{1}\text{a}(\xi(z))\,\text{b}(\xi(z))\;\left[\text{W}(\xi(z))\;\partial_{z}\xi(z)\right]\;dz\;.$ (35)
In Eq. (35), we have used the fact that $d\xi=\partial_{z}\xi(z)\,dz$ to write the integral explicitly in $z$. According to Eq. (34), there may exist coordinate parameters, $\text{x}_{0}$ and $\text{x}_{2}$, such that $\xi(z=0)=0$, and $\xi(z=1)=1$, which would mean that the limits of integration do not change.
@@
$\quad\quad\times\;\left(1-(\text{x}_{2}+1)z^{\text{x}_{0}}\right)^{-\text{B}_{0}-\text{B}_{1}-2}\;.$
While Eq. (34) provides a coordinate that modifies the boundary behavior of $\text{W}(\xi)$, it is Eq. (36) that must be analyzed to determine the special values of $\text{x}_{0}$ and $x_{2}$ that prevent the effective weight function from diverging at the domain boundaries. In essence, we have two independent conditions: (i) $\text{W}_{\text{eff.}}(0)=0$, and (ii) $\text{W}_{\text{eff.}}(1)=0$. We also have two unknowns: $\text{x}_{0}$ and $\text{x}_{1}$. Thus, $\text{x}_{0}$ and $\text{x}_{1}$ may be determined (up to positive constants) by careful analysis of $\text{W}_{\text{eff.}}$.
This analysis is somewhat tedious, but tractable. First, one should focus on the factor of $z^{\text{x}_{0}(\text{B}_{0}+1)-1}$ within $\text{W}_{\text{eff.}}$. If $\text{x}_{0}$ is chosen appropriately, then the real part of $\text{x}_{0}(\text{B}_{0}+1)-1$ should equal some positive constant, $\text{K}_{0}$, thereby allowing $\text{W}_{\text{eff.}}$ to go to zero as $z$ goes to zero. Since, $\text{B}_{0}$ is complex valued, $\text{x}_{0}$ must be also. These ideas result in the following constraints,
$\text{K}_{0}$ $=-1+\left[1+\text{Re}(\text{B}_{0})\right]\text{Re}(\text{x}_{0})-\text{Im}(\text{B}_{0})\text{Im}(\text{x}_{0}),$ (37a)
@@
If we assume that $\text{Re}(\text{x}_{0})$ only depends on $\text{Re}(\text{B}_{0})$, then then the result for $\text{x}_{0}$ follows,
@@
Given, $\text{x}_{0}$, one may then determine $\text{x}_{2}$. Returning to $\text{W}_{\text{eff.}}$, if its exponential factor has a negative argument as $z\rightarrow 1$, then $\text{W}_{\text{eff.}}$ is assured to be regular in that limit. To that end, one should consider $e^{\frac{\text{B}_{2}\text{x}_{2}}{z^{\text{x}_{0}}-1}}$, and seek $\text{x}_{2}$ such that
@@
In Eq. (41), argmin outputs the values for which $S$ is minimized in $\text{K}_{0}$ and $\text{K}_{2}$, $\text{mean}_{z}$ computes the unweighted mean over values of $z$, $\text{min}_{z}$ returns the smallest value of a quantity in $z$, and $\text{max}_{z}$ returns the largest value of a quantity in $z$.
In practice, one has a number of remaining options for actually evaluating Eq. (35). For the evaluation of Eq. (41), it is found to be useful to densely sample, on log-scale, $z$ near 0 and $1$. The coordinate freedom associated with $\text{x}_{0}$ introduces oscillations into $\xi(z)$ which must be well resolved for the accurate determination of $\text{K}_{0}$ and $\text{K}_{2}$. Similarly, for the evaluation of Eq. (35), one will want to densely sample near the boundaries; it is found that $z=\frac{1}{2}(1+\sin(t\pi/2))$, with $t$ uniformly spaced on (1,1)(-1,1), is sufficient for this purpose. For the determination of $\text{K}_{0}$ and $\text{K}_{2}$, and for the evaluation of Eq. (35), it is found that at least $10^{4}$ points in $z$ are needed for robust results. With these details in hand, it is now possible to compute the integration path, $\xi(z)$, and related scalar-products, Eq. (35), using e.g. $5^{\text{th}}$ order spline integration Virtanen et al. (2020).